Steam enters a 10,000 kW turbine of an ideal Rankine cycle at 10 Bar and 360 degree Celsius and expands to 0.1 bar before entering a condenser. Determine

A. Mass Flow Rate(kg/hr)

B. Condenser heat transfer(kW)

C. Net Power(kW)

D. Efficiency
Answer #1

First of all list out all the properties form the steam table at given conditions

As in the question it is given that turbines capacity to produce power is 10000kw so we first calculate work done per unit mass and then to the mass flow rate.

After that for net power output pump work is also required which then is calculated .

Mass flow rate = 14.458kg/s

=0.2409 kg/hr

Please refer images for detailed solution.

1 (360°c) 3 B=0.1 Whimbine = 10,000 kw Boiler foressure = lo bar (360°c) = P1 Enden som fressure = 0.1 bar = Pz from Steam ta2 = 6583-0.649 7.502 X2=0.7909 M = haces the hfg (P2) | he = 191,8 +0.7409 x 2392-9. ha = 2084.5577 RJ l May now, Wterbineshow, net power an be given as, Wnet = Wesbine-Wpomp. Wpump = hfq-hf (ps) 3 - Veces (P1-Ps) = 0.001010(m? 1 kq) x (10-0. Uxlo3. Q = hi-hfa =1776.2 - 192.799. P1= 2583.401 19/kg Layde = 690.64 kb/ta 2583.401 13/kg | Muy dle = 0.267 26.76